OpenStudy (anonymous):

urn contains 4 white and 6 red rolls. Four balls are drawn at random (without replacement) from the urn. Find the probability distribution of number of white balls?

terenzreignz (terenzreignz):

This follows a hypergeometric distribution. Its probability density function is given by \[\huge \frac{\left(\begin{matrix}N \\ n\end{matrix}\right)\left(\begin{matrix}K-N \\ k-n\end{matrix}\right)}{\left(\begin{matrix}K \\ k\end{matrix}\right)}\] Where N is the number of favourable outcomes, n is how many such draws are needed, and K is the total number of outcomes.

OpenStudy (anonymous):

how to solve it

terenzreignz (terenzreignz):

What is your favourable outcome? Drawing a white ball, right?

OpenStudy (anonymous):

4/10

terenzreignz (terenzreignz):

No, just 4. There are 4 white balls, so N = 4.

OpenStudy (anonymous):

oh yes, i calculate probability of white balss

OpenStudy (anonymous):

can any body solve this in simple form so that i can understand

OpenStudy (perl):

we can do it in steps

OpenStudy (perl):